# Representation Learning for Distributional Perturbation Extrapolation

### v1

@@ Abstract
We consider the problem of modelling the effects of unseen perturbations such as gene knockdowns or drug combinations on low-level measurements such as RNA sequencing data. Specifically, given data collected under some perturbations, we aim to predict the distribution of measurements for new perturbations. To address this challenging extrapolation task, we posit that perturbations act additively in a suitable, unknown embedding space. More precisely, we formulate the generative process underlying the observed data as a latent variable model, in which perturbations amount to mean shifts in latent space and can be combined additively. Unlike previous work, we prove that, given sufficiently diverse training perturbations, the representation and perturbation effects are identifiable up to affine transformation, and use this to characterize the class of unseen perturbations for which we obtain extrapolation guarantees. To estimate the model from data, we propose a new method, the perturbation distribution autoencoder (PDAE), which is trained by maximising the distributional similarity between true and predicted perturbation distributions. The trained model can then be used to predict previously unseen perturbation distributions. Empirical evidence suggests that PDAE compares favourably to existing methods and baselines at predicting the effects of unseen perturbations.
@@ 1 Introduction
Due to technological progress, large-scale perturbation data is becoming more abundant across several scientific fields. This is particularly the case for single-cell biology, where advancements in gene editing, sequencing, and mass spectrometry have led to the collection of vast transcriptomic or proteomic databases for various drug and gene perturbations (Wang et al., 2009; Jinek et al., 2012; Weinstein et al., 2013; Dixit et al., 2016; Norman et al., 2019). However, the exponential number of possible combinations of perturbations renders exhaustive experimentation prohibitive. Observations are thus typically only available for a subset of perturbations of interest, e.g., some single and double gene knockdowns or certain dosages of drugs. Having models capable of extrapolating to unseen combinations of perturbations, e.g., new multi-gene knockdowns or dosage combinations, would be highly desirable.
Prior Work.
@@
Several recent works advocate for the use of machine learning for biological perturbation modelling (Lopez et al., 2018, 2023), often with the goal of generalizing to new cell types (Lotfollahi et al., 2019; Bunne et al., 2023), unseen combinations of perturbations (Lotfollahi et al., 2023; Bereket & Karaletsos, 2023), or entirely new perturbations by leveraging the molecular structure of the involved compounds (Yu & Welch, 2022; Hetzel et al., 2022; Qi et al., 2024) or prior knowledge about gene-gene interactions (Roohani et al., 2024; Kamimoto et al., 2023). A common theme is the use of nonlinear representation learning techniques such as autoencoders (Rumelhart et al., 1986; Hinton & Salakhutdinov, 2006; Kingma & Welling, 2014) to embed observations in a latent space, in which the effects of perturbations are assumed to take on a simpler (e.g., additive) form. However, existing studies are purely empirical and lack theoretical underpinning. Despite promising results, the capabilities and fundamental limitations of existing methods thus remain poorly understood.
Overview and Contributions.
@@
In this work, we present a principled, theoretically-grounded approach for perturbation extrapolation. Given the unpaired nature of the available non-i.i.d. data (there are no correspondences between observations from different domains since each cell is only measured under one experimental condition) we consider the task of predicting population-level effects of perturbations, which we formalize as a distributional regression problem (§ 2). We then postulate a generative model (§ 3) which, similar to prior works, assumes that perturbations act as mean shifts in a suitable latent space, see Fig. 1 for an overview. For the case of a deterministic decoder, we analyse this model class theoretically (§ 4), proving that, under mild assumptions, the latent representation and the relative training perturbation effects are identifiable up to affine transformation if the available environments are sufficiently diverse (Thm. 4.1). This result implies extrapolation guarantees for unseen perturbations that can be expressed as linear combinations of training perturbations (Thm. 4.6). Based on these insights, we devise an autoencoder-based estimation method (§ 5) which draws inspiration from distributional principal autoencoders (Shen & Meinshausen, 2024a) and uses the energy score (Gneiting & Raftery, 2007) to assess the distributional similarity between predicted and ground-truth perturbation data. In simulations (§ 7), our approach compares favourably to the compositional perturbation autoencoder (CPA; Lotfollahi et al., 2023) and several baselines in terms of mean prediction and distributional fit.
Notation.
@@
(a) Simulated Multi-Domain Perturbation Data
[…]
(b) Graphical Model Representation
Figure 1: Task Description and Example Following the Assumed Data Generating Process. (a) During training, we are given $M\!=\!5$ training data sets in observation space (right, contour plots in grey), each of which is generated under a known combination of $K\!=\!3$ elementary perturbations. The corresponding (training) perturbation labels $\mathbf{l}_{e}$ are shown below the plots. During testing, we are given a new perturbation label and the task is to predict the corresponding distribution in observation space (right, blue and orange). We tackle this task by assuming that the effect of perturbations is linear additive in a suitable latent space (left). Both plots show kernel density estimates of the distributions. (b) For each experiment or environment $e\in[M]_{0}$, the corresponding dataset comprises a perturbation label $\bm{l}_{e}$ and $N_{e}$ observations $\bm{x}_{e,i}$. Perturbations are assumed to act as mean-shifts on a latent basal state, $\mathbf{z}^{\text{pert}}_{e,i}=\mathbf{z}^{\text{base}}_{e,i}+\mathbf{W}\mathbf{l}_{e}$ for an unknown perturbation matrix $\mathbf{W}$. An (unknown) stochastic nonlinear decoder with noise $\bm{\varepsilon}_{e,i}$ then yields the observed $\mathbf{x}_{e,i}=\mathbf{f}(\mathbf{z}^{\mathrm{pert}}_{e,i},\bm{\varepsilon}_{e,i})$. In the example in (a), $\mathbf{f}$ is constant in its second argument. Shaded and white nodes indicate observed and unobserved/latent variables, respectively.
2 Problem Setting: Distributional Perturbation Extrapolation
Observables.
Let $\mathbf{x}\in\mathbb{R}^{{d_{X}}}$ be an observation (e.g., omics data) that is obtained under one of several possible experimental conditions, and let $\mathbf{X}$ be the corresponding random vector. We model these conditions as combinations of $K$ elementary perturbations, each of which we assume can be encoded by a real number. Further, let $\mathbf{l}\in\mathbb{R}^{K}$ be a vector of perturbation labels that indicates if, or how much of, each perturbation was applied before collecting $\mathbf{x}$.
Example 2.1 (Gene perturbations).
For data arising from gene knockouts, a perturbation can be represented by a binary $\mathbf{l}\in\{0,1\}^{K}$, where $K$ denotes the number of potential targets and $l_{k}=1$ if and only if target $k$ was subject to a knockout experiment. For example, $\mathbf{l}=(1,0,1,0,0)$ indicates a multi-gene knockout on targets one and three.
Example 2.2 (Drug perturbations).
For data arising from applying varying amounts of $K$ different drugs, perturbations can be represented by continuous, non-negative labels $\mathbf{l}\in\mathbb{R}_{+}^{K}$, where $l_{k}$ indicates the (relative or absolute) amount of drug $k$ that was administered.
Data.
We have access to $M+1$ experimental datasets $\mathcal{D}_{0},\mathcal{D}_{1},...,\mathcal{D}_{M}$, each comprising a sample of $N_{e}$ observations $\mathbf{x}$ and a perturbation label $\mathbf{l}$, i.e., for all experiments or environments $e\in[M]_{0}$,
$\mathcal{D}_{e}=\left(\left(\mathbf{x}_{e,i}\right)_{i=1}^{N_{e}},\mathbf{l}_{e}\right).$ (2.1)
For all $e\in[M]_{0}$, we assume that $\mathbf{x}_{e,1},...,\mathbf{x}_{e,N_{e}}$ are i.i.d. realization from some underlying distribution $\mathbb{P}_{\mathbf{X}|\mathbf{l}_{e}}$.
Task.
Given data in the form of (2.1), we aim to predict the effects of new perturbations $\mathbf{l}_{\mathrm{test}}\not\in\{\mathbf{l}_{0},\mathbf{l}_{1},...,\mathbf{l}_{M}\}$ without observing any data from this condition (“zero-shot”). In particular, we are interested in the distribution over observations resulting from $\mathbf{l}_{\mathrm{test}}$. That is, we aim to leverage the training domains (2.1) to learn a map
$\mathbf{l}\mapsto\mathbb{P}_{\mathbf{X}|\mathbf{l}},$ (2.2)
which extrapolates beyond the training support of $\mathbf{l}$, i.e., the predictions should remain reliable for new inputs $\mathbf{l}_{\mathrm{test}}$.
Distributional vs. Mean Prediction.
Since (2.2) targets the full conditional distribution—rather than, say, the conditional mean $\mathbb{E}[\mathbf{X}|\mathbf{l}]$—it constitutes a (multi-variate) distributional regression task (Koenker & Bassett Jr, 1978; Koenker, 2005), also referred to as probabilistic forecasting (Gneiting & Raftery, 2007) or conditional generative modelling (Mirza, 2014; Sohn et al., 2015; Winkler et al., 2019). Therefore, we refer to the problem setting addressed in this paper as distributional perturbation extrapolation.
On Extrapolation.
Formally, extrapolation means that the value of the function (2.2) at $\mathbf{l}_{\mathrm{test}}$ is determined by its values on the training support $\{\mathbf{l}_{0},\mathbf{l}_{1},...,\mathbf{l}_{M}\}$. Intuitively, for this to be feasible without imposing very restrictive assumptions on the form of the mapping in (2.2), $\mathbf{l}_{\mathrm{test}}$ should be somehow related to the training perturbations $\mathbf{l}_{e}$. For example, given data resulting from individual perturbations, predict the effects of combinations thereof. This type of extrapolation to new combinations of inputs is also called compositional generalization (Lake et al., 2017; Goyal & Bengio, 2022). It is known to be challenging (Schott et al., 2022; Montero et al., 2021, 2022) and requires assumptions that sufficiently constrain the model class  (Wiedemer et al., 2024a, b; Lachapelle et al., 2023; Brady et al., 2023, 2025; Dong & Ma, 2022; Lippl & Stachenfeld, 2024).
3 Model: Perturbations as Mean Shifts in Latent Space
We now specify a generative process for the observed data in (2.1). In so doing, we aim to strike a balance between imposing sufficient structure on (2.2) to facilitate extrapolation, while remaining flexible enough to model the complicated, nonlinear effects which perturbations may have on the distribution of observations. Similar to prior work (Lotfollahi et al., 2023; Bereket & Karaletsos, 2023), we model the effect of perturbations as mean shifts in a latent space with perturbation-relevant latent variables $\mathbf{z}\in\mathbb{R}^{{d_{Z}}}$, [^1] Xu et al. (2024) propose tests for this assumption, given suitable data from individual and combinations of perturbations. which are related to the observations $\mathbf{x}$ via a nonlinear (stochastic) mixing function or generator $\mathbf{f}$. The full generative process amounts to a hierarchical latent variable model, which additionally contains noise variables ${\bm{\varepsilon}}$ that capture other variation underlying the observations $\mathbf{x}$, and which is represented as a graphical model in Fig. 1(b). Specifically, we posit for all $e\in[M]_{0}\cup\{\mathrm{test}\}$ and all $i\in[N_{e}]$:
$\mathbf{Z}^{\mathrm{base}}_{e,i}$ $\sim\mathbb{P}_{\mathbf{Z}},$ (3.1)
$\mathbf{Z}^{\mathrm{pert}}_{e,i}$ $:=\mathbf{Z}^{\mathrm{base}}_{e,i}+\mathbf{W}\mathbf{l}_{e},$ (3.2)
${\bm{\varepsilon}}_{e,i}$ $\sim\mathbb{Q}_{\bm{\varepsilon}},$ (3.3)
$\mathbf{X}_{e,i}$ $:=\mathbf{f}\left(\mathbf{Z}^{\mathrm{pert}}_{e,i},{\bm{\varepsilon}}_{e,i}\right),$ (3.4)
where $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are independent and identically distributed (i.i.d.) according to $\mathbb{P}_{\mathbf{Z}}$, and $({\bm{\varepsilon}}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are i.i.d. according to $\mathbb{Q}_{\bm{\varepsilon}}$ and jointly independent of $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$.
The basal state $\mathbf{Z}^{\mathrm{base}}$ in (3.1) describes the unperturbed state of latent variables, which can, in principle, be affected by perturbations, and is distributed according to a base distribution $\mathbb{P}_{\mathbf{Z}}$. The perturbation matrix $\mathbf{W}\in\mathbb{R}^{{d_{Z}}\times K}$ in (3.2) captures the effect of the $K$ elementary perturbations encoded in $\mathbf{l}$ on the latents and turns $\mathbf{Z}^{\mathrm{base}}$ into perturbed latents $\mathbf{Z}^{\mathrm{pert}}$. Since the same perturbation $\mathbf{l}_{e}$ is applied for all $i\in[N_{e}]$, all intra-dataset variability in $\mathbf{Z}^{\mathrm{pert}}$ is due to $\mathbb{P}_{\mathbf{Z}}$. The noise variables ${\bm{\varepsilon}}\in\mathbb{R}^{{d_{\epsilon}}}$ in (3.3) capture all other variation in the observed data that is unaffected by perturbations. W.l.o.g., we assume that it is distributed according to a fixed, uninformative distribution $\mathbb{Q}_{\bm{\varepsilon}}$ such as a standard isotropic Gaussian. The noise serves as an additional input to the (stochastic) mixing function or generator $\mathbf{f}:\mathbb{R}^{d_{Z}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{d_{X}}$ in (3.4), which produces observations for the perturbed latent. This generative model allows us to capture any conditional distribution $\mathbb{P}_{\mathbf{X}|\mathbf{Z}=\mathbf{z}}$ (see, e.g., Peters et al., 2017, Prop. 7.1) and is more flexible than relying on parametric assumptions, such as a Gaussian (Lotfollahi et al., 2023) or negative Binomial (Bereket & Karaletsos, 2023; Lopez et al., 2018, 2023) likelihood parametrised by a deterministic decoder.
For a given $e$ and $\mathbf{l}_{e}$, the generative process in (3.1)–(3.4) induces a distribution over observations $\mathbf{x}$, which we denote by $\mathbb{P}_{e}$ or $\mathbb{P}_{\mathbf{X}|\mathbf{l}_{e}}$, defined as the push-forward of $\mathbb{P}_{\mathbf{Z}}$ and $\mathbb{Q}_{\bm{\varepsilon}}$ through (3.2) and (3.4), such that:
[…]
4 Theory: Identifiability and Extrapolation Guarantees
In this section, we present our theoretical analysis for the model class introduced in § 3. Here, we assume that the decoder is deterministic, so that we can write $\mathbf{f}:\mathbb{R}^{d_{Z}}\to\mathbb{R}^{d_{X}}$, see Remark 4.5 below for further discussion.
4.1 Identifiability
We first study identifiability, that is, the question under what assumptions and up to what ambiguities certain parts of the postulated generative process can be provably recovered assuming access to the full distributions. As established by the following results, our model class is identifiable up to affine transformation, provided that the training perturbations are sufficiently diverse, the dimension of the latent space is known and some additional technical assumptions hold.
Theorem 4.1 (Affine identifiability for Gaussian latents).
For $M\in\mathbb{Z}_{\geq 0}$, let $\mathbf{l}_{0},...,\mathbf{l}_{M}\in\mathbb{R}^{K}$ be ${M{+}1}$ perturbation labels. Let $\mathbf{f},\widetilde{\mathbf{f}}:\mathbb{R}^{{d_{Z}}}\to\mathbb{R}^{d_{X}}$, $\mathbf{W},\widetilde{\mathbf{W}}\in\mathbb{R}^{{d_{Z}}\times K}$, and $\mathbb{P},\widetilde{\mathbb{P}}$ be distributions on $\mathbb{R}^{{d_{Z}}}$ such that the models $(\mathbf{f},\mathbf{W},\mathbb{P})$ and $(\widetilde{\mathbf{f}},\widetilde{\mathbf{W}},\widetilde{\mathbb{P}})$ induce the same observed distributions, i.e.,
[…]
Assume further that:

### v2

# Representation Learning for Extrapolation in Perturbation Modeling

@@ Abstract
We consider the problem of modeling the effects of perturbations, such as gene knockdowns or drugs, on measurements, such as single-cell RNA or protein counts. Given data for some perturbations, we aim to predict the distribution of measurements for new combinations of perturbations. To address this challenging extrapolation task, we posit that perturbations act additively in a suitable, unknown embedding space. We formulate the data-generating process as a latent variable model, in which perturbations amount to mean shifts in latent space and can be combined additively. We then prove that, given sufficiently diverse training perturbations, the representation and perturbation effects are identifiable up to orthogonal transformation and use this to characterize the class of unseen perturbations for which we obtain extrapolation guarantees. We establish a link between our model class and shift interventions in linear latent causal models. To estimate the model from data, we propose a new method, the perturbation distribution autoencoder (PDAE), which is trained by maximizing the distributional similarity between true and simulated perturbation distributions. The trained model can then be used to predict previously unseen perturbation distributions. Through simulations, we demonstrate that PDAE can accurately predict the effects of unseen but identifiable perturbations, supporting our theoretical results.
Machine Learning, ICML
@@ 1 Introduction
Due to technological progress, large-scale perturbation data is becoming increasingly abundant across several scientific fields. In single-cell biology, for example, advancements in gene editing, sequencing, and mass spectrometry have led to the collection of vast transcriptomic or proteomic databases for various drug and gene perturbations (Wang et al., 2009; Jinek et al., 2012; Dixit et al., 2016; Norman et al., 2019; Zhang et al., 2025). However, the exponential number of possible combinations of perturbations renders exhaustive experimentation prohibitive. Observations are thus typically only available for a subset of perturbations of interest, e.g., some single and double gene knockdowns or certain dosages of drugs. Having models capable of extrapolating to unseen combinations of perturbations, e.g., new multi-gene knockdowns or dosage combinations, would be highly desirable.
@@
Prior work. Several studies advocate for the use of machine learning for biological perturbation modeling (Lopez et al., 2018, 2023; Hugi et al., 2025), often aiming to generalize to new cell types (Lotfollahi et al., 2019; Bunne et al., 2023; Adduri et al., 2025); unseen combinations of perturbations (Lotfollahi et al., 2023; Bereket & Karaletsos, 2023); entirely new perturbations by leveraging the molecular structure of the involved compounds (Yu & Welch, 2022; Hetzel et al., 2022; Qi et al., 2024) or known gene-gene interactions (Roohani et al., 2024; Kamimoto et al., 2023); or combinations thereof (He et al., 2025; Adduri et al., 2025). A common theme is the use of nonlinear representation learning techniques such as autoencoders to embed observations in a latent space, in which the effects of perturbations are assumed to take on a simpler (e.g., additive) form. However, existing studies are purely empirical and lack theoretical underpinning. The capabilities and fundamental limitations of existing methods thus remain poorly understood.
(a) Simulated multi-domain perturbation data
$\mathbf{Z}^{\mathrm{base}}_{e,i}$$\mathbf{Z}^{\mathrm{pert}}_{e,i}$$\mathbf{a}_{e}$${\bm{\varepsilon}}_{e,i}$$\mathbf{X}_{e,i}$ $N_{e}$ $M+1$
(b) Graphical model representation
Figure 1: Task description and example following the assumed data generating process. (a) During training, we are given M=5M\!=\!5 training data sets in observation space (right, contour plots in grey), each of which is generated under a known combination of K=3K\!=\!3 elementary perturbations. The corresponding (training) perturbation labels $\mathbf{a}_{e}$ are shown below the plots. During testing, we are given a new perturbation label and the task is to predict the corresponding distribution in observation space (right, blue and orange). We tackle this task by assuming that the effect of perturbations is linear additive in a suitable latent space (left). Both plots show kernel density estimates of the distributions. (b) For each experiment or environment $e\in[M]_{0}$, the corresponding dataset comprises a perturbation label $\mathbf{a}_{e}$ and $N_{e}$ observations $\bm{x}_{e,i}$. Perturbations are assumed to act as mean-shifts on a latent basal state, $\mathbf{z}^{\text{pert}}_{e,i}=\mathbf{z}^{\text{base}}_{e,i}+\mathbf{W}\mathbf{a}_{e}$ for an unknown perturbation matrix $\mathbf{W}$. An (unknown) stochastic nonlinear decoder with noise $\bm{\varepsilon}_{e,i}$ then yields the observed $\mathbf{x}_{e,i}=\mathbf{f}(\mathbf{z}^{\mathrm{pert}}_{e,i},\bm{\varepsilon}_{e,i})$. In the example in (a), $\mathbf{f}$ is constant in its second argument. Shaded and white nodes indicate observed and unobserved/latent variables, respectively.
@@
Overview and contributions. In this work, we present theoretically-grounded perspective on perturbation extrapolation. Given the typically unpaired nature of the available non-i.i.d. data—e.g., each cell can only be measured once since the measurement process is destructive—we consider the task of predicting population-level perturbation effects, which we formalize as a distributional regression problem (§˜2). We then postulate a generative model (§˜3) which, similar to prior works, assumes that perturbations act as mean shifts in a suitable latent space, see Fig.˜1 for an overview. For the case of a deterministic decoder, we analyse this model class theoretically (§˜4), proving that, under mild assumptions, the latent representation and the relative training perturbation effects are identifiable up to orthogonal transformation if the available environments are sufficiently diverse (Thm.˜4.1). This result implies extrapolation guarantees for certain unseen perturbations that can be expressed as linear combinations of training perturbations (Thm.˜4.7). By considering shift interventions in linear structural equation models, we establish a connection between our model class and causal representation learning (Prop.˜4.9). Based on our theoretical results, we then propose an autoencoder-based estimation method (§˜5) which uses the energy score (Gneiting & Raftery, 2007) to assess distributional similarity between predicted and ground-truth perturbation data. In simulations (§˜7), this approach compares favourably to relevant baselines in terms of mean prediction and distributional fit.
Notation. We write scalars as $a$, column-vectors as $\mathbf{a}$, and matrices as $\mathbf{A}$. We use uppercase for random variables and lowercase for their realizations. The pushforward of a distribution $\mathbb{P}$ by a measurable function $f$ is denoted by $f_{\#}\mathbb{P}$. The Euclidean ($\ell_{2}$) norm is denoted by $\left\lVert\cdot\right\rVert$. For $n\in\mathbb{N}_{\geq 1}$, we use the shorthands $n]:=\{1,...,n\}$ and $n]_{0}:=[n]\cup\{0\}$.
2 Distributional perturbation extrapolation
Observables. Let $\mathbf{x}\in\mathbb{R}^{{d_{X}}}$ be an observation (e.g., omics data) that is obtained under one of several possible experimental conditions, and let $\mathbf{X}$ be the corresponding random vector. We model these conditions as combinations of $K$ elementary perturbations, each of which we assume can be encoded by a real number. Further, let $\mathbf{a}\in\mathbb{R}^{K}$ be a vector of perturbation labels that indicates if, or how much of, each perturbation was applied before collecting $\mathbf{x}$.
Example 2.1 (Gene perturbations).
For data arising from gene knockouts, a perturbation can be represented by a binary $\mathbf{a}\in\{0,1\}^{K}$, where $K$ denotes the number of potential targets and $a_{k}=1$ if and only if target $k$ was subject to a knockout experiment. For example, $\mathbf{a}=(1,0,1,0,0)$ indicates a multi-gene knockout on targets one and three.
Example 2.2 (Drug perturbations).
For data arising from applying varying amounts of $K$ different drugs, perturbations can be represented by continuous, non-negative labels $\mathbf{a}\in\mathbb{R}_{+}^{K}$, where $a_{k}$ indicates the (relative or absolute) amount of drug $k$ that was administered.
Data. We have access to $M+1$ experimental datasets $\mathcal{D}_{0},\mathcal{D}_{1},...,\mathcal{D}_{M}$, each comprising a sample of $N_{e}$ observations $\mathbf{x}$ and a perturbation label $\mathbf{a}$, i.e., for all experiments or environments $e\in[M]_{0}$,
$\mathcal{D}_{e}=\big((\mathbf{x}_{e,i})_{i=1}^{N_{e}},\mathbf{a}_{e}\big).$ (2.1)
For all $e\in[M]_{0}$, we assume that $\mathbf{x}_{e,1},...,\mathbf{x}_{e,N_{e}}$ are i.i.d. realizations from some underlying distribution $\mathbb{P}_{\mathbf{X}|\mathbf{a}_{e}}$.
Task. Given data in the form of (2.1), we aim to predict the effects of new perturbations $\mathbf{a}_{\mathrm{test}}\not\in\{\mathbf{a}_{0},\mathbf{a}_{1},...,\mathbf{a}_{M}\}$ without observing any data from this condition (“zero-shot”). In particular, we are interested in the distribution over observations resulting from $\mathbf{a}_{\mathrm{test}}$. That is, we aim to leverage the training domains (2.1) to learn a map
$\mathbf{a}\mapsto\mathbb{P}_{\mathbf{X}|\mathbf{a}},$ (2.2)
which extrapolates beyond the training support of $\mathbf{a}$, i.e., the predictions should remain reliable for new inputs $\mathbf{a}_{\mathrm{test}}$.
Distributional vs mean prediction. Since (2.2) targets the full conditional distribution—rather than, say, the conditional mean $\mathbb{E}[\mathbf{X}|\mathbf{a}]$—it constitutes a (multi-variate) distributional regression task (Koenker & Bassett Jr, 1978), a.k.a. probabilistic forecasting (Gneiting & Raftery, 2007) or conditional generative modeling (Sohn et al., 2015). Therefore, we refer to the problem setting addressed in this paper as distributional perturbation extrapolation.
On extrapolation. Formally, extrapolation means that the value of the function (2.2) at $\mathbf{a}_{\mathrm{test}}$ is determined by its values on the training support $\{\mathbf{a}_{0},\mathbf{a}_{1},...,\mathbf{a}_{M}\}$. Intuitively, for this to be feasible without imposing very restrictive assumptions on the form of the mapping in (2.2), $\mathbf{a}_{\mathrm{test}}$ should be somehow related to the training perturbations $\mathbf{a}_{e}$. For example, given data resulting from individual perturbations, predict the effects of combinations thereof. This type of extrapolation to new combinations of inputs is also called compositional generalization (Lake et al., 2017; Goyal & Bengio, 2022). It is known to be challenging (Schott et al., 2022; Montero et al., 2021, 2022) and requires assumptions that sufficiently constrain the model class  (Wiedemer et al., 2024a, b; Lachapelle et al., 2023; Brady et al., 2023, 2025; Dong & Ma, 2022; Lippl & Stachenfeld, 2024).
3 Perturbations as mean shifts in latent space
We now specify a generative process for the observed data in (2.1). In so doing, we aim to strike a balance between imposing sufficient structure on (2.2) to facilitate extrapolation, while remaining flexible enough to model the complicated, nonlinear effects which perturbations may have on the distribution of observations. Similar to other approaches (e.g., Lotfollahi et al., 2023; Bereket & Karaletsos, 2023), we model the effect of perturbations as mean shifts in a latent space with perturbation-relevant latent variables $\mathbf{z}\in\mathbb{R}^{{d_{Z}}}$, which are related to the observations $\mathbf{x}$ via a nonlinear (stochastic) mixing function or generator $\mathbf{f}$. The full generative process amounts to a hierarchical latent variable model, which additionally contains noise variables $\bm{\varepsilon}$} that capture other variation underlying the observations $\mathbf{x}$, and which is represented as a graphical model in Fig.˜1(b). Specifically, we posit for all $e\in[M]_{0}\cup\{\mathrm{test}\}$ and all $i\in[N_{e}]$:
$\mathbf{Z}^{\mathrm{base}}_{e,i}$ $\sim\mathbb{P}_{\mathbf{Z}},\qquad$ $\mathbf{Z}^{\mathrm{pert}}_{e,i}$ $:=\mathbf{Z}^{\mathrm{base}}_{e,i}+\mathbf{W}\mathbf{a}_{e},$ (3.1)
${\bm{\varepsilon}}_{e,i}$ $\sim\mathbb{Q}_{\bm{\varepsilon}},\qquad$ $\mathbf{X}_{e,i}$ $:=\mathbf{f}\big(\mathbf{Z}^{\mathrm{pert}}_{e,i},{\bm{\varepsilon}}_{e,i}\big),$ (3.2)
where $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are i.i.d. according to $\mathbb{P}_{\mathbf{Z}}$, and $({\bm{\varepsilon}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$ are i.i.d. according to $\mathbb{Q}_{\bm{\varepsilon}}$ and jointly independent of $(\mathbf{Z}^{\mathrm{base}}_{e,i})_{e\in[M]_{0},i\in[N_{e}]}$.
The basal state $\mathbf{Z}^{\mathrm{base}}$ in (3.1) describes the unperturbed state of latent variables, which can, in principle, be affected by perturbations, and is distributed according to a base distribution $\mathbb{P}_{\mathbf{Z}}$. The perturbation matrix $\mathbf{W}\in\mathbb{R}^{{d_{Z}}\times K}$ captures the effect of the $K$ elementary perturbations encoded in $\mathbf{a}$ on the latents and turns $\mathbf{Z}^{\mathrm{base}}$ into perturbed latents $\mathbf{Z}^{\mathrm{pert}}$. Since the same perturbation $\mathbf{a}_{e}$ is applied for all $i\in[N_{e}]$, all within-dataset variability in $\mathbf{Z}^{\mathrm{pert}}$ is due to $\mathbb{P}_{\mathbf{Z}}$. The noise variables ${\bm{\varepsilon}}\in\mathbb{R}^{{d_{\epsilon}}}$ in (3.2) capture all other variation in the observed data that is unaffected by perturbations. W.l.o.g., we assume that it is distributed according to a fixed, uninformative distribution $\mathbb{Q}_{\bm{\varepsilon}}$ such as a standard isotropic Gaussian. The noise serves as an additional input to the (stochastic) mixing function or generator $\mathbf{f}:\mathbb{R}^{d_{Z}}\times\mathbb{R}^{d_{\epsilon}}\to\mathbb{R}^{d_{X}}$, which produces observations for the perturbed latent. This implicit generative model can capture any conditional distribution $\mathbb{P}_{\mathbf{X}|\mathbf{Z}=\mathbf{z}}$ (e.g., Peters et al., 2017, Prop. 7.1) and is more flexible than relying on parametric assumptions, such as a Gaussian (Lotfollahi et al., 2023) or negative Binomial (Bereket & Karaletsos, 2023; Lopez et al., 2018, 2023) likelihood parametrised by a deterministic decoder.
For a given $e$ and $\mathbf{a}_{e}$, the generative process in (3.1) and (3.2) induces a distribution over observations $\mathbf{x}$, which we denote by $\mathbb{P}_{e}$ or $\mathbb{P}_{\mathbf{X}|\mathbf{a}_{e}}$, defined as the push-forward of $\mathbb{P}_{\mathbf{Z}}$ and $\mathbb{Q}_{\bm{\varepsilon}}$ through (3.1) and (3.2), such that for all $e\in[M]_{0}$:
$(\mathbf{X}_{e,i})_{i\in[N_{e}]}\overset{\text{i.i.d.}}{\sim}\mathbb{P}_{e}\,.$ (3.3)
4 Theoretical results
We now present our theoretical analysis for the model class from §˜3. Here, we assume that the decoder is deterministic, so that we can write $\mathbf{f}:\mathbb{R}^{d_{Z}}\to\mathbb{R}^{d_{X}}$}, see Remark˜4.6 for further discussion. All proofs are provided in Appx.˜A.
4.1 Identifiability
We first study identifiability, i.e., under which assumptions and up to which ambiguities certain parts of the postulated generative process can be provably recovered given access to the full distributions. As established by the following results, our model class is identifiable up to orthogonal transformation, provided that the training perturbations are sufficiently diverse, the dimension of the latent space is known, and some additional technical assumptions hold.
Theorem 4.1 (Identifiability up to orthogonal transformation).
@@ 4.1 Identifiability
$\mathbf{f}\left(\mathbf{Z}+\mathbf{W}\mathbf{a}_{e}\right)\overset{d}{=}\widetilde{\mathbf{f}}\Big(\widetilde{\mathbf{Z}}+\widetilde{\mathbf{W}}\mathbf{a}_{e}\Big),$ (4.1)
with $\mathbf{Z}\sim\mathbb{P}$ and $\widetilde{\mathbf{Z}}\sim\widetilde{\mathbb{P}}$ independent. Assume further that:
@@
    [Gaussianity] $\mathbf{Z}$ and $\widetilde{\mathbf{Z}}$ are standard isotropic multivariate Gaussians with means chosen such that the latent distributions in domain $e=0$ are centered at the origin, i.e., $\mathbb{P}=\mathcal{N}\left(-\mathbf{W}\mathbf{a}_{0},\mathbf{I}\right)$ and $\widetilde{\mathbb{P}}=\mathcal{N}(-\widetilde{\mathbf{W}}\mathbf{a}_{0},\mathbf{I})$;
  3. (iii)
    [sufficient diversity] the matrix $\widetilde{\mathbf{W}}\mathbf{A}\in\mathbb{R}^{{d_{Z}}\times M}$, where $\mathbf{A}\in\mathbb{R}^{K\times M}$ is the matrix with columns $(\mathbf{a}_{1}-\mathbf{a}_{0})$, …, $(\mathbf{a}_{M}-\mathbf{a}_{0})$, has full row rank, i.e., $\mathrm{rank}(\widetilde{\mathbf{W}}\mathbf{A})={d_{Z}}$, or the same holds for $\mathbf{W}\mathbf{A}$.
Then the latent representation and the effects of the observed perturbation combinations relative to $\mathbf{a}_{0}$ (as captured by $\mathbf{W}\mathbf{A}$) are identifiable up to orthogonal transformation, i.e., there is an orthogonal matrix $\mathbf{O}\in O({d_{Z}})$ such that
$\forall\mathbf{z}\in\mathbb{R}^{d_{Z}}:\qquad\widetilde{\mathbf{f}}^{-1}\circ\mathbf{f}(\mathbf{z})$ $=\mathbf{O}\mathbf{z},$ (4.2)
$\widetilde{\mathbf{W}}\mathbf{A}$ $=\mathbf{O}\mathbf{W}\mathbf{A}.$ (4.3)
Corollary 4.2.
If, in addition to the assumptions of Thm.˜4.1, $\mathbf{A}\in\mathbb{R}^{K\times M}$ has full row rank (i.e., $\mathrm{rank}(\mathbf{A})=K\leq M$), then the perturbation matrix $\mathbf{W}$ is identifiable up to orthogonal transformation, i.e., $\widetilde{\mathbf{W}}=\mathbf{O}\mathbf{W}$, where $\mathbf{O}\in O({d_{Z}})$ is an orthogonal matrix.